Solve.
x^2+y^2=97
xy=36

Answer by richwmiller(17219)

factors of 36
12 and 3 squares 144 is too much
9 and 4 81 and 16 97 is just right
6 and 6 36 and 36 72
x=36/y
(36/y)^2+y^2=97
1296/y^2+y^2=97
y^2-13y+36 = 0
(y-9) (y-4) = 0
y=9 and y=4